# Crystal-field magnetostriction of the spin ice under ultrahigh magnetic fields

### v1

@@ Abstract
We present a comprehensive study of the magnetoelastic properties of the Ising pyrochlore oxide Ho2Ti2O7, known as spin ice, by means of high-field magnetostriction measurements and numerical calculations. When a magnetic field is applied along the crystallographic $\langle 111\rangle$ axis, the longitudinal magnetostriction exhibits a broad maximum in the low-field regime around 30 T, followed by a dramatic lattice contraction due to crystal-field (CF) level crossing at $B_{\rm cf}\sim 65$ T. The transverse magnetostriction exhibits a contrasting behavior, highlighting the anisotropic nature of the CF striction. We identify distinct timescales of spin dynamics and CF-phonon dynamics by applying a magnetic field with different field-sweep rates. Our mean-field calculations, based on a point-charge model, successfully reproduce the overall magnetostriction behavior, revealing the competition between the exchange striction and CF striction. A signature of the CF level crossing is also observed through adiabatic magnetocaloric-effect measurements, consistent with our magnetostriction data.
@@ I Introduction
There are several well-known microscopic origins of magnetostriction MDoerr_2005 . One such mechanism involves the dependence of an ionic radius on its electronic configuration. A field-induced change in the ionic state can result in significant volume magnetostriction, as exemplified by a spin-state transition in LaCoO3 AIkeda_2020 and valence transitions in $f$ electron-based valence-fluctuating materials KYoshimura_1988 ; Mus_2004 ; AMiyake_2022 . In contrast to these on-site mechanisms, exchange striction refers to the change in distance between two magnetic ions to minimize exchange coupling energy [Fig. 1(a)]. Here, $\Delta L/L$ reflects changes in local spin correlations $\langle{\bm{S}}_{i}\cdot{\bm{S}}_{j}\rangle$ between neighboring sites $i$ and $j$, and consequently correlates with the magnetization $M$ Zapf_2008 ; MJaime_2012 ; AIkeda_2019 ; AMiyata_2021 . The crystal field (CF) also influences the arrangement of surrounding anions and the wave function of the magnetic site to minimize the Coulomb energy. Applying a magnetic field alters the wave function through CF level hybridization, potentially causing anisotropic lattice deformation, known as CF striction [Fig. 1(b)]. Although magnetostriction is universally observed in all magnetic materials, its theoretical description is far from straightforward and is often simplified by phenomenological approaches Patri_2019 or effective magnetoelastic models Kimura_2014 ; AMiyata_2020 .
@@
The target compound in this study is Ho2Ti2O7, one of the most extensively studied $4f$ Ising pyrochlore magnet, known as a classical spin ice Harris_1997 ; ROsenkranz_2000 ; STBramwell_2001 ; Petrenko_2003 ; Fennell_2009 ; YNakanishi_2011 ; CCastelnovo_2012 ; SErfanifam_2014 ; MRuminyDFT_2016 ; MRuminy_2016_CF ; LOpherden_2019 ; Wang_2021 . The strong Ising anisotropy towards the $\langle 111\rangle$ axes arises from a combination of dominant nearest-neighbor ferromagnetic interactions and the CF effect. The strong geometrical frustration results in a highly degenerate ground state, characterized by short-range spin correlations where two spins point inward and two outward within each tetrahedron (“2-in–2-out”) Petrenko_2003 [Fig. 1(c)]. When a magnetic field is applied along the [111] direction, a metamagnetic transition occurs at approximately 2 T, flipping one spin and resulting in a “3-in-1-out” (or “1-in-3-out”) configuration in the tetrahedra, representing a fully polarized state within the Ising limit [Fig. 1(c)]. In this field orientation, each tetrahedron has one magnetic site with the easy axis parallel to $B$ (the “easy-axis (EA) site”) and three sites with axes not parallel to $B$ (the “hard-axis (HA) sites”).
@@
Figure 1: (a)(b) Schematics of (a) exchange striction and (b) cyrstal-field (CF) striction. (c) Three types of spin configurations when a magnetic field is applied along [111] on Ho2Ti2O7. The field orientation is the easy axis for the spin colored by orange (EA site), whereas the hard axis for the other three spins colored by green (HA site). “2-in, 2-out” represents the highly-degenerate spin-ice state, “3-in, 1-out” the excited state, and “all-up” the forced ferromagnetic state beyond the Ising limit. (d) Energy level diagram for the single-ion Hamiltonian of Eq. (2) when one of the three HA spins along the hard axis is subjected to a magnetic field along the [111] direction. The two lowest CF levels hybridize at $B_{\rm cf}^{\rm cal}=67$ T.
@@
Given the strong spin-orbit coupling intertwining spin and orbital degrees of freedom, the CF-phonon interaction is crucial for the magnetoelastic properties. However, as the easy-axis directions vary at the four magnetic sites in Ho2Ti2O7, local lattice deformations induced by the CF tend to cancel out, making its effect on the crystal structure less evident. To clarify this, the observation of the CF striction in high magnetic fields would be useful. In Ho2Ti2O7, the energy gap between the lowest and the first-excited CF levels is approximately 260 K in zero field MRuminy_2016_CF . The application of an external magnetic field along [111] can hybridize these CF levels at the HA sites, as shown in the CF level diagram in Fig. 1(d), eventually realizing a forced ferromagnetic (“all-up”) state [Fig. 1(c)]. The CF level crossing is theoretically expected at $B_{\rm cf}^{\rm cal}=67$ T based on the exact diagonalization of the CF Hamiltonian. Indeed, Erfanifam et al. observed a signature of the CF level hybridization for $B\parallel[111]$ through magnetization and ultrasound measurements up to 60 T SErfanifam_2014 , and Opherden et al. observed a sequence of metamagnetic transitions accompanied by the CF level crossing for $B\parallel[5513]$ up to 120 T LOpherden_2019 .
@@
Here, we investigate the CF striction in Ho2Ti2O7 using magnetostriction measurements under ultrahigh magnetic fields up to 120 T, complemented by magnetization and magnetocaloric-effect (MCE) measurements. We observe significant anisotropic CF striction, characterized by a lattice contraction amounting to $\Delta L/L\sim-5\times 10^{-4}$ along the field direction. Our mean-field calculations, based on a point-charge model and incorporating CF, two-ion exchange interactions, phonons, and CF-phonon coupling, successfully reproduce the magnetostriction data. Similar high-field magnetostriction is observed in another spin ice compound, Pr2Zr2O7, suggesting that the CF striction is a common feature in rare-earth-based spin-ice systems.
@@ II.1 Experiments
Single crystals of Ho2Ti2O7 and Pr2Zr2O7 were grown from polycrystalline feed rods using a high-temperature xenon-type optical floating zone furnace. Crystal orientations were checked using a backscattering Laue x-ray diffractometer. Several pieces of as-grown crystals were cut and polished into rectangular parallelepipeds with dimensions of approximately 1 $\times$ 1 $\times$ 3 mm3.
@@
  $M$ HSTC Nondestructive short pulse $\mu$s 4 ms 130 T 62 T Quasi-adiabatic Nonadiabatic
  $\Delta L/L$ HSTC Nondestructive long pulse $\mu$s 36 ms 120 T 55 T Quasi-adiabatic Nonadiabatic
  MCE  Nondestructive long pulse  36 ms    55 T   Quasi-adiabatic & Nonadiabatic
@@
In the nondestructive pulsed magnet, the magnetization and magnetostriction measurements were performed in either liquid-4He or gas 4He environment, i.e., nonadiabatic conditions. In the HSTC system, the sample was cooled to approximately 5 K using a liquid-4He flow-type cryostat, where the measurement condition should be quasi-adiabatic because of the short pulsed-field duration. To monitor the temperature change of the sample during the field sweep, we measured the magnetocaloric effect (MCE) up to 55 T under both nonadiabatic and quasi-adiabatic conditions using the nondestructive long-pulsed magnet 2013_Kih . A sensitive Au16Ge84 film thermometer was sputtered on the surface of the crystal, with temperature calibration performed using a commercial RuO2 thermometer.
@@ II.2 Theoretical calculations
where the sum is taken over all nearest-neighbor pairs, $J(\bm{R}_{ij})\equiv J_{\rm dipo}+J_{\rm ex}$ is the distance-dependent effective exchange coupling, and $\bm{R}_{ij}$ is the vector connecting sites $i$ and $j$. Assuming that $J$ is linearly modulated by a slight bond-length change $\epsilon$, we expand $J({\bm{R}}_{ij})$ as,
$J(\bm{R}_{ij})\approx J(\bm{R}_{ij}^{0})+\sum_{\alpha,\beta,\gamma}\frac{\partial J}{\partial R^{\alpha}}\frac{\partial\epsilon_{\alpha\gamma}R_{ij}^{\gamma}}{\partial\epsilon_{\beta}}\epsilon_{\beta},$ (4)
where $\bm{R}_{ij}^{0}$ is the distance between the original positions of sites $i$ and $j$, and the indices $\alpha,\beta,\gamma=x,y,z$ refer to an Euclidean coordinate system parallel to the crystal axes, i.e., $x\parallel a$, $y\parallel b$ and $z\parallel c$. For the lattice contribution, we consider a phonon term, $\hat{H}_{\rm ph}$, and a CF-phonon term, $\hat{H}_{\rm cfph}$, which represents the coupling between the CF and phonons. The detailed formulas for these two terms are provided in Appendix B. Combining all these terms, we construct the total Hamiltonian as
@@
We performed numerical simulations on Eq. (5) using the McPhase program MRotter_2002 ; MRotter_2004 ; MRotter_2012 , which specializes in calculating the physical properties of $4f$-based magnets within the framework of mean-field theory MRotter_2007 ; JJensen_2007 ; TStoter_2020 . McPhase employs a point-charge model to determine the CF wavefunctions and CF energy gap notes . To match the experimental energy gap of approximately 260 K ROsenkranz_2000 ; MRuminy_2016_CF , we apply a scaling factor of 0.7 to the nominal charges Ho ($3+$), Zr ($4-$), O ($2-$). To set the initial state in zero field, we utilized the crystallographic parameters of Ho2Ti2O7 reported in Ref. Baroudi_2015 . We set the dipolar interaction to $J_{\rm dipo}=2.4$ K STBramwell_2001 and the nearest-neighbor exchange interaction to $J_{\rm ex}=0.1$ K, yielding a total effective exchange coupling of $J=2.5$ K. This set of parameters successfully reproduces both the metamagnetic transition in the low-field regime and the CF level crossing at $B_{\rm cf}\sim 65$ T, as experimentally observed. We note that even a small change of 0.1 K in $J_{\rm ex}$ results in a 6 T shift in $B_{\rm cf}^{\rm cal}$. In Eq. (4), we assume $dJ/dR\simeq dJ_{\rm ex}/dR$, which is determined by the Bethe-Slater curve, as shown in Fig. 8 in Appendix B.
@@
where $Z=\sum_{n}\exp(-E_{n}/k_{\rm B}T)$ is the partition function for sublattice $i$, and $|n\rangle$ is the eigenstate corresponding to the eigenvalue $E_{n}$ of $\hat{H}_{\rm single-ion}+\hat{H}_{\rm ex}$. The total magnetization $M$ is obtained by averaging the magnetization of all sublattices, in the unit of $\mu_{\rm B}$/Ho.
For the magnetostriction, we calculate strains self-consistently based on Eq. (5) and the following equation:
@@
$+\frac{1}{2}\sum_{i,i^{\prime},\delta,\delta^{\prime},\alpha^{\prime},\gamma}\frac{\partial J_{\delta\delta^{\prime}}(\bm{R}_{ii^{\prime}})}{\partial R^{\alpha^{\prime}}}\frac{\partial\epsilon_{\alpha^{\prime}\gamma}\partial R^{\gamma}_{ii^{\prime}}}{\partial\epsilon_{\alpha}}\langle\hat{J}_{i\delta}\hat{J}_{i^{\prime}\delta^{\prime}}\rangle,$
where $i$ and $i^{\prime}$ denote atomic site positions, while $\delta$, $\delta^{\prime}$, and $\alpha^{\prime}$, $\gamma$ run over the spatial coordinates $x$, $y$, and $z$, represented as the integers 1, 2, and 3, respectively. $c^{\alpha\beta}$ denotes the elastic constant, where $\alpha$ and $\beta$ are indices in Voigt notaion, ranging from 1 to 6. $G_{\rm mix}^{\alpha\delta}$ represents the phonon-strain coupling, and $G_{\rm cfph}^{\alpha lm}$ represents the CF-phonon coupling. For details of these parameters, see Appendix B. In Eq. (7), the first and second terms on the right side contribute to the CF striction, and the third term is the exchange striction. The total lattice strain, i.e., magnetostriction, can be obtained by
$\frac{\Delta L}{L}=\sum_{ij}\epsilon_{ij}{\bm{{l}}_{i}}{\bm{l}_{j}},$ (8)
where $\bm{l}$ denotes the unit vector in the direction of measurement. Strain $\epsilon_{ij}$ is a $3\times 3$ symmetric tensor, hence has 6 independent matrix elements: $\epsilon_{xx}$, $\epsilon_{yy}$, $\epsilon_{zz}$, $\epsilon_{xy}$, $\epsilon_{xz}$, $\epsilon_{yz}$. We obtain the longitudinal magnetostriction ($\Delta L/L\parallel$ [111]) by calculating $\Delta L/L=(\epsilon_{xx}+\epsilon_{yy}+\epsilon_{zz}+2(\epsilon_{xy}+\epsilon_{xz}+\epsilon_{yz}))/3$. In this study, we use Voigt notation, where $\epsilon_{ij}$ is relabeled as $\epsilon_{1}=\epsilon_{xx}$, $\epsilon_{2}=\epsilon_{yy}$, $\epsilon_{3}=\epsilon_{zz}$, $\epsilon_{4}=2\epsilon_{yz}=2\epsilon_{zy}$, $\epsilon_{5}=2\epsilon_{xz}=2\epsilon_{zx}$, and $\epsilon_{6}=2\epsilon_{xy}=2\epsilon_{yx}$.

### v3

# Crystal field magnetostriction of spin ice under ultrahigh magnetic fields

@@ Abstract
We present a comprehensive study of the magnetoelastic properties of the Ising pyrochlore oxide Ho2Ti2O7, known as spin ice, by means of high-field magnetostriction measurements and numerical calculations. When a magnetic field is applied along the crystallographic $\langle 111\rangle$ axis, the longitudinal magnetostriction exhibits a broad maximum in the low-field regime around 30 T, followed by a dramatic lattice contraction due to crystal-field (CF) level crossing at $B_{\rm cf}\sim 65$ T. The transverse magnetostriction exhibits a contrasting behavior, highlighting the anisotropic nature of the CF striction. By applying a magnetic field at varying sweep rates, we identify distinct timescales of spin dynamics that are relevant to monopole formation and annihilation, as well as CF-phonon dynamics. Our mean-field calculations, based on a point-charge model, successfully reproduce the overall magnetostriction behavior, revealing the competition between the exchange striction and CF striction. A signature of the CF level crossing is also observed through adiabatic magnetocaloric-effect measurements, consistent with our magnetostriction data.
@@ I Introduction
There are several well-known microscopic origins of magnetostriction MDoerr_2005 . One such mechanism involves the dependence of an ionic radius on its electronic configuration. A field-induced change in the ionic state can result in significant volume magnetostriction, as exemplified by a spin-state transition in LaCoO3 AIkeda_2020 and valence transitions in $f$ electron-based valence-fluctuating materials KYoshimura_1988 ; Mus_2004 ; AMiyake_2022 . In contrast to these on-site mechanisms, exchange striction refers to the change in distance between two magnetic ions to minimize exchange coupling energy [Fig. 1(a)]. Here, $\Delta L/L$ reflects changes in local spin correlations $\langle{\bm{S}}_{i}\cdot{\bm{S}}_{j}\rangle$ between neighboring sites $i$ and $j$, and consequently correlates with the magnetization $M$ Zapf_2008 ; MJaime_2012 ; AIkeda_2019 ; AMiyata_2021 . The crystal field (CF) also influences the arrangement of surrounding anions and the wavefunction of the magnetic site to minimize Coulomb energy. Applying a magnetic field alters the wavefunction through CF level hybridization, potentially causing anisotropic lattice deformation, known as CF striction [Fig. 1(b)]. Although magnetostriction is universally observed in all magnetic materials, its theoretical description is far from straightforward and is often simplified by phenomenological approaches Patri_2019 or effective magnetoelastic models Kimura_2014 ; AMiyata_2020 .
@@
The target compound in this study is Ho2Ti2O7, one of the most extensively studied $4f$ Ising pyrochlore magnet, known as a classical spin ice Harris_1997 ; ROsenkranz_2000 ; KMatsuhira_2000 ; STBramwell_2001 ; Petrenko_2003 ; Fennell_2009 ; YNakanishi_2011 ; CCastelnovo_2012 ; SErfanifam_2014 ; MRuminyDFT_2016 ; MRuminy_2016_CF ; LOpherden_2019 ; Wang_2021 . The strong Ising anisotropy along the $\langle 111\rangle$ axes arises from a combination of dominant nearest-neighbor ferromagnetic interactions and the CF effects. This leads to a highly degenerate ground state characterized by short-range spin correlations, where two spins point inward and two point outward within each tetrahedron (“2-in–2-out”) Petrenko_2003 [Fig. 1(c)]. When a magnetic field is applied along the [111] direction, a metamagnetic transition occurs at approximately 2 T, flipping one spin and resulting in a “3-in–1-out” (or “1-in–3-out”) configuration in the tetrahedra, representing a fully polarized state within the Ising limit [Fig. 1(c)]. In this field orientation, each tetrahedron has one magnetic site with the easy axis parallel to $B$ [the easy-axis (EA) site] and three sites with axes not parallel to $B$ [the hard-axis (HA) sites].
@@
Figure 1: (a), (b) Schematics of (a) exchange striction and (b) crystal-field (CF) striction. The orange arrows indicate the magnetic moments, and the yellow (blue) spheres represent the valence electron density of the magnetic ion (ligands). The elliptical deformation of the yellow spheres schematically illustrates the crystal-field anisotropy, which aligns with the magnetic moment direction. The thick black arrows indicate the direction of ion displacement when the magnetic moment orientation is altered by an external magnetic field $B$. (c) Three types of spin configurations when a magnetic field is applied along [111] on Ho2Ti2O7. The spins aligned along the easy axis are colored orange (EA site), while those aligned along the hard axis are colored green (HA site). (d) Energy level diagram for the single-ion Hamiltonian of Eq. (2), where one of the three HA spins is subjected to a magnetic field in the [111] direction. The two lowest CF levels hybridize at $B_{\rm cf}^{\rm cal}=67$ T.
@@
Given the strong spin-orbit coupling that links spin and orbital degrees of freedom, the CF-phonon interaction plays a key role in magnetoelastic properties, forming the foundation of magnetostrictive phenomena. However, as the easy-axis directions vary at the four magnetic sites in Ho2Ti2O7, local lattice deformations induced by the CF tend to cancel out, making its effect on the crystal structure less evident. To clarify this, the observation of the CF striction in high magnetic fields would be useful. In Ho2Ti2O7, the energy gap between the lowest and the first-excited CF levels is approximately 260 K in zero field ROsenkranz_2000 ; MRuminy_2016_CF . The application of an external magnetic field along [111] can hybridize these CF levels at the HA sites, as shown in the CF level diagram in Fig. 1(d), eventually realizing a forced ferromagnetic (all-up) state [Fig. 1(c)]. The CF level crossing is theoretically expected at $B_{\rm cf}^{\rm cal}=67$ T based on the exact diagonalization of the CF Hamiltonian. Indeed, Erfanifam et al. observed a signature of the CF level hybridization for $B\parallel[111]$ through magnetization and ultrasound measurements up to 60 T SErfanifam_2014 , and Opherden et al. observed a sequence of metamagnetic transitions accompanied by the CF level crossing for $B\parallel[5513]$ up to 120 T LOpherden_2019 .
@@
Here, we investigate the CF striction in Ho2Ti2O7 using magnetostriction measurements under ultrahigh magnetic fields up to 120 T, complemented by magnetization and magnetocaloric effect (MCE) measurements. We observe significant anisotropic CF striction, characterized by a lattice contraction amounting to $\Delta L/L\sim-5\times 10^{-4}$ along the field direction. We were also able to resolve the timescale of spin dynamics, finding it to be comparable to the microsecond range, relevant to monopole formation and annihilation. Additionally, CF-phonon dynamics appear to occur at rates that are comparable to, or potentially even faster than, the microsecond range. Our mean-field calculations, based on a point-charge model and incorporating CF, two-ion exchange interactions, phonons, and CF-phonon coupling, successfully reproduce the magnetostriction data. Similar high-field magnetostriction is observed in another spin ice compound, Pr2Zr2O7, suggesting that the CF striction is a common feature in rare-earth-based spin-ice systems.
@@ II.1 Experiments
Single crystals of Ho2Ti2O7 and Pr2Zr2O7 were grown from polycrystalline feed rods using a high-temperature xenon-type optical floating zone furnace at ISSP, University of Tokyo. Crystal orientations were checked using a backscattering Laue x-ray diffractometer. Several pieces of as-grown crystals were cut and polished into rectangular parallelepipeds with dimensions of approximately 1 $\times$ 1 $\times$ 3 mm3.
@@
  $M$ HSTC Nondestructive short pulse $\mu$s 4 ms 130 T 62 T Quasiadiabatic Nonadiabatic
  $\Delta L/L$ HSTC Nondestructive long pulse $\mu$s 36 ms 120 T 55 T Quasiadiabatic Nonadiabatic
  MCE  Nondestructive long pulse  36 ms    55 T   Quasiadiabatic & nonadiabatic
@@
In the nondestructive pulsed magnet, the magnetization and magnetostriction measurements were performed in either liquid-4He or gas 4He environment, i.e., nonadiabatic conditions. In the HSTC system, the sample was cooled to approximately 5 K using a liquid-4He flow-type cryostat, where the measurement condition should be quasiadiabatic because of the short pulsed-field duration. To monitor the temperature change of the sample during the field sweep, we measured the magnetocaloric effect (MCE) up to 55 T under both nonadiabatic and quasiadiabatic conditions using the nondestructive long-pulsed magnet 2013_Kih . A sensitive Au16Ge84 film thermometer was sputtered on the surface of the crystal, with temperature calibration performed using a commercial RuO2 thermometer.
@@ II.2 Theoretical calculations
where the sum is taken over all nearest-neighbor pairs, $J(\bm{R}_{ij})\equiv J_{\rm dipo}+J_{\rm ex}$ is the distance-dependent effective exchange coupling, and $\bm{R}_{ij}$ is the vector connecting sites $i$ and $j$. Assuming that $J$ is linearly modulated by a small strain $\epsilon$, we expand $J({\bm{R}}_{ij})$ as,
$J(\bm{R}_{ij})\approx J(\bm{R}_{ij}^{0})+\sum_{\alpha,\gamma=1,2,3,\beta=1,...,6}\frac{\partial J}{\partial R^{\alpha}}\frac{\partial\epsilon_{\alpha\gamma}R_{ij}^{\gamma}}{\partial\epsilon_{\beta}}\epsilon_{\beta},$ (4)
where $\bm{R}_{ij}^{0}$ is the distance between the original positions of sites $i$ and $j$, and the indices $\alpha,\gamma=x,y,z$ refer to an Euclidean coordinate system parallel to the crystal axes, i.e., $x\parallel a$, $y\parallel b$ and $z\parallel c$. The index $\beta$ follows Voigt notation, ranging from 1 - 6. In other words, $\epsilon_{\beta}$ represents the strain tensor $\epsilon_{\alpha\gamma}$ in Voigt notation. For the lattice contribution, we consider a phonon term, $\hat{H}_{\rm ph}$, and a CF-phonon term, $\hat{H}_{\rm cfph}$, which represents the coupling between the CF and phonons. The detailed formulas for these two terms are provided in Appendix B. Combining all these terms, we construct the total Hamiltonian as
@@
We performed numerical simulations based on Eq. (5) using the McPhase program MRotter_2002 ; MRotter_2004 ; MRotter_2012 , which specializes in calculating the physical properties of $4f$-based magnets within the framework of mean-field theory MRotter_2007 ; JJensen_2007 ; TStoter_2020 . McPhase employs a point-charge model to determine the CF wavefunctions and CF energy gap notes . To match the experimental energy gap of approximately 260 K ROsenkranz_2000 ; MRuminy_2016_CF , we apply a scaling factor of 0.7 to the nominal charges Ho ($3+$), Ti ($4+$), O ($2-$). To set the initial state in zero field, we utilized the crystallographic parameters of Ho2Ti2O7 reported in Ref. Baroudi_2015 . We set the dipolar interaction to $J_{\rm dipo}=2.4$ K STBramwell_2001 and the nearest-neighbor exchange interaction to $J_{\rm ex}=0.1$ K, yielding a total effective exchange coupling of $J=2.5$ K. This set of parameters successfully reproduces both the metamagnetic transition in the low-field regime and the CF level crossing at $B_{\rm cf}\sim 65$ T, as experimentally observed. We note that even a small change of 0.1 K in $J_{\rm ex}$ results in a 6 T shift in $B_{\rm cf}^{\rm cal}$. In Eq. (4), we assume $dJ/dR\simeq dJ_{\rm ex}/dR$, which is determined by the Bethe-Slater curve, as shown in Fig. 8 in Appendix B.
@@
where $Z=\sum_{n}\exp(-E_{n}/k_{\rm B}T)$ is the partition function for sublattice $i$, and $|n\rangle$ is the eigenstate corresponding to the eigenvalue $E_{n}$ of $\hat{H}_{\rm total}$ treated in mean field approximation. The total magnetization $M$ is obtained by averaging the magnetization of all sublattices, in the unit of $\mu_{\rm B}$/Ho.
For the magnetostriction, we calculate strains self-consistently using Eq. (5) and the following equation:
@@
$+\frac{1}{2}\sum_{i,i^{\prime},\delta,\delta^{\prime},\alpha^{\prime},\gamma=1,2,3}\frac{\partial J_{\delta\delta^{\prime}}(\bm{R}_{ii^{\prime}})}{\partial R^{\alpha^{\prime}}}\frac{\partial\epsilon_{\alpha^{\prime}\gamma}\partial R^{\gamma}_{ii^{\prime}}}{\partial\epsilon_{\alpha}}\langle\hat{J}_{i\delta}\hat{J}_{i^{\prime}\delta^{\prime}}\rangle,$
where $i$, $i^{\prime}$, $\delta$, $\delta^{\prime}$, $\alpha^{\prime}$, and $\gamma$ run over the spatial coordinates $x$, $y$, and $z$, represented as the integers 1, 2, and 3, respectively. $c^{\alpha\beta}$ denotes the elastic constant, where $\alpha$ and $\beta$ are indices in Voigt notaion, ranging from 1 - 6. $G_{\rm mix}^{\alpha\delta}$ represents the phonon-strain coupling, and $G_{\rm cfph}^{\alpha lm}$ represents the CF-phonon coupling. For details of these parameters, see Appendix B. In Eq. (7), the first and second terms on the right side contribute to the CF striction, and the third term is the exchange striction. The total lattice strain, i.e., magnetostriction, can be obtained by
$\frac{\Delta L}{L}=\sum_{i,j=1,2,3}\epsilon_{ij}{\bm{{l}}_{i}}{\bm{l}_{j}},$ (8)
where $\bm{l}$ denotes the unit vector in the direction of measurement. Strain $\epsilon_{ij}$ is a $3\times 3$ symmetric tensor, hence has six independent matrix elements: $\epsilon_{xx}$, $\epsilon_{yy}$, $\epsilon_{zz}$, $\epsilon_{xy}$, $\epsilon_{xz}$, $\epsilon_{yz}$. We obtain the longitudinal magnetostriction ($\Delta L/L\parallel$ [111]) by calculating $\Delta L/L=(\epsilon_{xx}+\epsilon_{yy}+\epsilon_{zz}+2(\epsilon_{xy}+\epsilon_{xz}+\epsilon_{yz}))/3$. In this study, we use Voigt notation, where $\epsilon_{ij}$ is relabeled as $\epsilon_{1}=\epsilon_{xx}$, $\epsilon_{2}=\epsilon_{yy}$, $\epsilon_{3}=\epsilon_{zz}$, $\epsilon_{4}=2\epsilon_{yz}=2\epsilon_{zy}$, $\epsilon_{5}=2\epsilon_{xz}=2\epsilon_{zx}$, and $\epsilon_{6}=2\epsilon_{xy}=2\epsilon_{yx}$.